RodeWolf:

Between 1990 and 1999, the number of movie screens in the United States increased by about 1500 each year. In 1996, there were 29,690 movie screens. Write an equation of a line, in slope-intercept form, to find the total number of screens y for any year x.

jimthompson5910:

How far did you get with this?

RodeWolf:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 How far did you get with this? \(\color{#0cbb34}{\text{End of Quote}}\) NO where i am stumped on this

jimthompson5910:

The portion that says "increased by about 1500 each year" indicates the slope is m = 1500 This is because the slope represents the rate of change. Agreed or no?

RodeWolf:

Ok I agree

jimthompson5910:

Also, the info that "In 1996, there were 29,690 movie screens" indicates x = 6 and y = 29690 pair up together. The x value represents the number of years after 1990. Using x = 6, y = 29690, and m = 1500, we can solve for b in the equation below y = mx+b 29690 = 1500*6+b What do you get when you solve for b?

RodeWolf:

ok give me a sec let me work that out

RodeWolf:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 Also, the info that "In 1996, there were 29,690 movie screens" indicates x = 6 and y = 29690 pair up together. The x value represents the number of years after 1990. Using x = 6, y = 29690, and m = 1500, we can solve for b in the equation below y = mx+b 29690 = 1500*6+b What do you get when you solve for b? \(\color{#0cbb34}{\text{End of Quote}}\) where did 6 come from?

jimthompson5910:

1996 is 6 years after 1990, so x = 6 represents this

RodeWolf:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 1996 is 6 years after 1990, so x = 6 represents this \(\color{#0cbb34}{\text{End of Quote}}\) oh ok thxs

jimthompson5910:

no problem

RodeWolf:

i'm almost done with solving

jimthompson5910:

alright

RodeWolf:

I don't think this is right. Ok I did 1500 divide 29690 which I got 0.050 then I brought the 6+b which I made 6b, so I divided 6 and 0.050 then got the answer b=0.0083

jimthompson5910:

What do you get when you evaluate 1500*6

RodeWolf:

wait ok yeah I didn't work that out right hold on i'm sorry thank you for helping me out

jimthompson5910:

no worries

RodeWolf:

Ok now I got 9000 when I did 6*1500

jimthompson5910:

correct, so we go from 29690 = 1500*6+b to 29690 = 9000+b This is the same as saying 29690 = b+9000

jimthompson5910:

How can we undo that "plus 9000"?

RodeWolf:

we would subtract 9000 from 29690 which then equals 20690 = b right?

jimthompson5910:

yep so because b = 20690 we go from y = mx+b to y = 1500x+20690 The y intercept 20690 is the number of screens in 1990. This is because x = 0 corresponds to the year 1990.

RodeWolf:

Ok that makes since now. Thank you so much jimthompson5910

jimthompson5910:

You're welcome